# Robust Estimation of Polyserial Correlation

### v1

@@ 7.3 Additional simulations
The third additional simulation, described in Appendix D.3, is concerned with distributional misspecification, where the polyserial model is misspecified for all observations in a sample (cf. Section 4.2). The simulation shows that the potential of robustness gain with our robust estimator depends on the case-specific characteristics of the nonnormal sampling distribution of $(X,\eta)$ for which the polyserial model is distributionally misspecified. In brief, if the sampling distribution can be reasonably well approximated by a mixture of a normal distribution and some other unknown distribution to emulate the contamination model in (11), then the robust estimator can improve upon the MLE. If the sampling distribution does not admit such an approximation, the robust estimator may not yield an improvement.
@@ 8 Empirical application
The ordinal variable epilie is the score of a lie scale and measures social desirability of a participant’s responses: The higher the score, the less trustworthy their responses. Originally, this variable has 10 unique ordinal response options. [^8] The Eysenck Personality Inventory constructs the variable epilie as the number of times that a given participant has responded “yes” in nine separate yes-no questions. Since higher scores are associated with higher levels of socially desirable responding, this variable can be interpreted as being ordinal. However, in the data of Revelle, (2025), no participant has the highest score of 9 or lowest score of 0, and only two participants have the second-highest score of 8. To avoid numerical instability in the estimation of thresholds, we therefore assign the two participants with score 8 the score 7, that is, we merge the seventh and eighth response category, and treat the score 1 as the first response category. Thus, the variable epilie here takes values in $\mathcal{Y}=\{1,2,\dots,7\}$ with a total of $r=7$ response options. Figure 3(b) visualizes the frequency of each response option.
@@ 9 Discussion and conclusion
As a remedy, we propose a novel estimator designed to be robust against partial misspecification of the polyserial model. The robust estimator leverages the theory of minimum density power divergence (DPD; Basu et al.,, 1998) estimation, which, as we show, is able to cope with mixed-type variables, unlike standard approaches for robust estimation with exclusively continuous or exclusively categorical variables. The proposed minimum DPD estimator achieves robustness by implicitly downweighting observations that cannot be sufficiently well fitted by the polyserial model. The ensuing weights are a useful analytical tool for identifying observations discrepant to the polyserial model, such as (but not limited to) outliers and careless responses. Our proposed methodology is implemented as part of the free open-source R package robcat (Welz et al., 2025a, ), which is publicly available at https://CRAN.R-project.org/package=robcat.
@@
This paper suggests a number of extensions. For instance, while our software implementation is reasonably fast with about 5 seconds for computing a robust estimate, substantial speedups could be achieved by rewriting the source code in C++ via Rcpp (Eddelbuettel,, 2013). Moreover, the ability to robustly estimate polyserial correlation could be particularly useful in structural equation modeling with mixed data to robustify such analyses against partial model misspecification. We leave these avenues to further research. Overall, we believe that our robust estimator can contribute to the growing literature on making latent variable analyses less dependent on easily violated modeling assumptions.
Appendix A Details on asymptotic properties
A.1 Alternative expressions of polyserial model density
@@ A.3 Main theorem
A.4 Uncertainty quantification: Comparison with ML
@@
A.5 Covariance matrix estimation
@@
Regarding computation, for a given $\bm{\theta}\in\bm{\Theta}$, computation of the objects $\big(\bm{B}\left(\bm{\theta}\right),\bm{\xi}\left({\bm{\theta}}\right),\bm{K}\left(\bm{\theta}\right)\big)$ is cheap because they only require computation of finite summations of terms with closed-form expressions. The only computational bottleneck is the matrix $\bm{A}\left(\bm{\theta}\right)$, which is a component of $\widehat{\bm{J}}_{N}\left(\bm{\theta}\right)$ and is defined as
@@
A.6 Gradient and Hessian expressions
@@
A.6.1 First order expressions
@@ A.6.2 Ancillary derivatives
A.6.3 Second order expressions
By the product rule, we obtain for the Hessian of ${p}_{XY}\left(x,y;\bm{\theta}\right)$ the expression
$\frac{\partial^{2}{p}_{XY}\left(x,y;\bm{\theta}\right)}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}=$
$\quad{p}_{X}\left(x;\bm{\theta}\right)\frac{\partial^{2}{p}_{Y|X}\left(y\ |\ x;\bm{\theta}\right)}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}+2\frac{\partial{p}_{X}\left(x;\bm{\theta}\right)}{\partial\bm{\theta}}\left(\frac{\partial{p}_{Y|X}\left(y\ |\ x;\bm{\theta}\right)}{\partial\bm{\theta}}\right)^{\top}+{p}_{Y|X}\left(y\ |\ x;\bm{\theta}\right)\frac{\partial^{2}{p}_{X}\left(x;\bm{\theta}\right)}{\partial{\bm{\theta}}\partial{\bm{\theta}}^{\top}}.$
We have established in Section A.6.1 expressions for the gradient of the marginal and conditional density. The (symmetric) Hessian matrix of the marginal density is given by
@@ A.6.3 Second order expressions
$\qquad\frac{1}{(1-\rho^{2})^{3/2}}\Bigg[\phi^{\prime}\left(\tau^{*}_{y}\left({x;\bm{\theta}}\right)\right)\left(\rho\tau_{y}-\frac{x-\mu}{\sigma}\right)\frac{\partial\tau^{*}_{y}\left({x;\bm{\theta}}\right)}{\partial\rho}+\tau_{y}\phi\left(\tau^{*}_{y}\left({x;\bm{\theta}}\right)\right)+$
$\qquad\qquad\qquad\qquad\phi^{\prime}\left(\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)\right)\left(\rho\tau_{y-1}-\frac{x-\mu}{\sigma}\right)\frac{\partial\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)}{\partial\rho}+\tau_{y-1}\phi\left(\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)\right)\Bigg]$
@@
$\frac{\partial^{2}{p}_{Y|X}\left(y\ |\ x;\bm{\theta}\right)}{\partial{\rho}\partial{\mu}}$ $=\frac{1}{\sigma\sqrt{1-\rho^{2}}}\Bigg[\bigg(1+\frac{\rho}{1-\rho^{2}}\bigg)\bigg(\phi\left(\tau^{*}_{y}\left({x;\bm{\theta}}\right)\right)-\phi\left(\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)\right)\bigg)+$
@@
$\frac{\partial^{2}{p}_{Y|X}\left(y\ |\ x;\bm{\theta}\right)}{\partial{\rho}\partial{(\sigma^{2})}}$ $=\frac{x-\mu}{\sigma^{3}\sqrt{1-\rho^{2}}}\Bigg[\bigg(1+\frac{\rho}{1-\rho^{2}}\bigg)\bigg(\phi\left(\tau^{*}_{y}\left({x;\bm{\theta}}\right)\right)-\phi\left(\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)\right)\bigg)+$
@@
for $k\in\{1,\dots,r-1\}$.
@@ Appendix B Algorithm for upper bound of raw weights
    Repeat Step 1 for all $y\in\mathcal{Y}$ and return the upper bound $S_{\alpha}(\bm{\theta})=\max\left\{S^{*}_{y}:y\in\mathcal{Y}\right\}$.
@@ C.3 Additional simulation results
Figure C.3: Fraction of the 5,000 simulation repetitions in which a point estimator did not numerically converge (top), and its associated standard error could not be computed due to singularity of $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator (bottom). The point estimators always converged, resulting in overplotted lines.
@@
Furthermore, Figure C.3 illustrates the fraction of the 5,000 repetitions in which the correlation estimate $\widehat{\rho}_{N}$ of given estimator $\widehat{\bm{\theta}}_{N}$ did not numerically converge to a solution. In similar fashion, it visualizes the fraction in which its associated estimated standard error $\widehat{\textnormal{SE}}\left(\widehat{\rho}_{N}\right)$ could not be computed due to singularity of matrix $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator in Section 6.2. The point estimates $\widehat{\rho}_{N}$ always numerically converged to a solution across all configurations and repetitions. Up to (and including) contamination fraction $\varepsilon=0.3$, the standard errors could virtually always be computed. Only in extremely high-contamination settings with $\varepsilon>0.3$, as the estimators gradually break down (Figure C.2), standard errors could occasionally also not be computed. In particular, at the highest considered contamination fraction $\varepsilon=0.49$, the estimator with tuning constant $\alpha=1$ fails to compute standard errors in about half of the repetitions. However, this estimator itself seems to be unstable in this setting to begin with (see Figure 2), which may cause frequent singularity of matrix $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$.

### v2

@@ 7.3 Additional simulations
The third additional simulation, described in Appendix D.3, is concerned with distributional misspecification, where the polyserial model is misspecified for all observations in a sample (cf. Section 4.2). The simulation shows that the potential of robustness gain with our robust estimator depends on the case-specific characteristics of the nonnormal sampling distribution of $(X,\eta)$ for which the polyserial model is distributionally misspecified. In brief, if the sampling distribution can be reasonably well approximated by a mixture of a normal distribution and some other unknown distribution to emulate the contamination model in (8), then the robust estimator can improve upon the MLE. If the sampling distribution does not admit such an approximation, the robust estimator may not yield an improvement but produce similar estimates as ML, which can be quite poor under distributional misspecification (e.g., Bedrick,, 1995, , and our simulations in Appendix D.3). Thus, we advice applied researchers to always test for partially-latent normality when ML and robust estimator yield similar estimates. While such similarity may be due to normality indeed holding true, it may also be due to distributional misspecification from a specific nonnormal distribution for which the robust estimator cannot improve over ML. We provide details and more explicit guidance in Appendix D.3.
@@ 8 Empirical application
The ordinal variable epilie is the score of a lie scale and measures social desirability of a participant’s responses: The higher the score, the less trustworthy their responses. Originally, this variable has 10 unique ordinal response options. [^13] The Eysenck Personality Inventory constructs the variable epilie as the number of times that a given participant has responded “yes” in nine separate yes-no questions. Since higher scores are associated with higher levels of socially desirable responding, this variable can be interpreted as being ordinal. However, in the data of Revelle, (2025), no participant has the highest score of 9 or lowest score of 0, and only two participants have the second-highest score of 8. To avoid numerical instability in the estimation of thresholds, [^14] It may happen that a robust estimator attempts to effectively eliminate thresholds corresponding to (nearly) empty cells in a certain way that we describe in detail in Appendix D.3.2, thereby causing instability. This phenomenon has been noted before by Welz et al., (2026). we therefore assign the two participants with score 8 the score 7, that is, we merge the seventh and eighth response category, and treat the score 1 as the first response category. Thus, the variable epilie here takes values in $\mathcal{Y}=\{1,2,\dots,7\}$ with a total of $r=7$ response options. Figure 3(b) visualizes the frequency of each response option.
@@ 9 Discussion and conclusion
As a remedy, we propose a novel estimator designed to be robust against partial misspecification of the polyserial model. The robust estimator leverages the theory of minimum density power divergence (DPD; Basu et al.,, 1998) estimation, which, as we show, is able to cope with mixed-type variables, unlike standard approaches for robust estimation with exclusively continuous or exclusively categorical variables. The proposed minimum DPD estimator achieves robustness by implicitly downweighting observations that cannot be sufficiently well fitted by the polyserial model. The ensuing weights are a useful analytical tool for identifying observations discrepant to the polyserial model, such as (but not limited to) outliers and careless responses. Our proposed methodology is implemented as part of the free open-source R package robcat (Welz et al.,, 2025), which is publicly available at https://CRAN.R-project.org/package=robcat.
@@
A central practical consideration is the choice of tuning constant $\alpha$, governing the robustness-efficiency tradeoff. While simulations suggest that $\alpha=0.5$ is a good compromise with about 76% of ML efficiency, we acknowledge that a detailed investigation is an important avenue of further research. We recommend to always compare various choices of $\alpha$, like in Figure 5, to assess estimation stability and severity of (partial) model misspecification. Alternatively, one may fix the efficiency loss and select the corresponding value of $\alpha$ as part of an ex-ante power analysis (cf. Basu et al.,, 1998, Section 5).
This paper suggests a number of extensions. For instance, while our software implementation is reasonably fast with about two seconds for computing a robust estimate, substantial speedups could be achieved by rewriting the source code in C++ via Rcpp (Eddelbuettel,, 2013). Moreover, the ability to robustly estimate polyserial correlation could be particularly useful in structural equation modeling with mixed data to robustify such analyses against partial model misspecification. We leave these avenues to further research. Overall, we believe that our robust estimator can contribute to the growing literature on making latent variable analyses less dependent on easily violated modeling assumptions.
Appendix A Technical details
This appendix section contains technical details on the polyserial model and robust estimation thereof. Concerning the polyserial model, Subsection A.1 covers identification of point polyserial correlation and Subsection A.2 provides alternative but equivalent definitions of the polyserial model density. Concerning estimation, Subsection A.3 introduces compact notation for the asymptotic analysis, Subsection A.4 rigorously establishes the main theorem, Subsection A.5 compares the asymptotic covariance matrices of ML to the robust estimator, Subsection A.6 discusses covariance matrix estimation, and Subsection A.7 provides closed-form gradient and Hessian expressions.
A.1 Identification of point polyserial correlation
Point polyserial correlation is the correlation between the observed variable $X$ and the observed ordinal variable $Y$ under the partially-latent normality assumption. To identify point polyserial correlation, one needs to introduce a scoring system for the response options of the ordinal $Y$. Recall that we assume that $Y$ takes values in the set $\mathcal{Y}=\{1,2,\dots,r\}$, where the elements in $\mathcal{Y}$ need not admit a numerical interpretation. Assigning a scoring system introduces a numeric meaning to each response category. We refer to Fernández et al., (2020), Ivanova & Berger, (2001), and Olsson et al., (1982, pp. 340–341) for discussions of different scoring systems.
$\begin{split}\mu_{Y}&=\mathbb{E}\left[Y\right]=\sum_{j=1}^{r}y_{j}\cdot{p}_{Y}\left(y_{j};\bm{\theta}\right),\qquad\textnormal{and}\\ \sigma^{2}_{Y}&=\mathbb{V}\mathrm{ar}\left[Y\right]=\sum_{j=1}^{r}y_{j}^{2}\cdot{p}_{Y}\left(y_{j};\bm{\theta}\right)-\mu_{Y}^{2}.\end{split}$ (A.1)
$\mu_{XY}=\mathbb{E}\left[XY\right]=\mu_{X}\left(y_{r}-\sum_{j=1}^{r-1}\Phi\left(\tau_{j}\right)\big(y_{j+1}-y_{j}\big)\right)+\rho\sigma_{X}\sum_{j=1}^{r-1}\phi\left(\tau_{j}\right)\big(y_{j+1}-y_{j}\big),$ (A.2)
$\widetilde{\rho}=\mathrm{Cor}\left[X,\ Y\right]=\frac{\mathrm{Cov}\left[X,\ Y\right]}{\sigma_{X}\sigma_{Y}}=\frac{1}{\sigma_{X}\sigma_{Y}}\left(\mu_{XY}-\mu_{X}\mu_{Y}\right),$ (A.3)
where explicit expressions for $(\mu_{Y},\sigma^{2}_{Y})$ and $\mu_{XY}$ are provided in Equations (A.1) and (A.2), respectively. The correlation coefficient $\widetilde{\rho}$ between the continuous $X$ and the ordinal $Y$ is referred to as point polyserial coefficient. For a given scoring system, $\widetilde{\rho}$ can be computed from the parameter vector $\bm{\theta}=\left(\rho,\mu,\sigma^{2},\bm{\tau}^{\top}\right)^{\top}$ of the polyserial model, a property established by Olsson et al., (1982). In particular, an estimator of $\widetilde{\rho}$ can be constructed by substituting $\bm{\theta}$ for an estimate thereof in Equations (A.1)–(A.3).
For the remainder of this appendix, we resort back to the original assumption that the support of the ordinal $Y$ is given by a set of adjacent integers without numeric interpretation, $\mathcal{Y}=\{1,2,\dots,r\}$, which, we reiterate, is without loss of generality for identifying the polyserial correlation coefficient.
A.2 Alternative expressions of the polyserial model density
@@ A.4 Main theorem
A.5 Uncertainty quantification: Comparison with ML
@@
A.6 Covariance matrix estimation
@@
Regarding computation, for a given $\bm{\theta}\in\bm{\Theta}$, computation of the objects $\big({\color[rgb]{0,0,0}\widehat{\bm{B}}_{N}\left(\bm{\theta}\right)},{\color[rgb]{0,0,0}\widehat{\bm{\xi}}_{N}\left(\bm{\theta}\right)},{\color[rgb]{0,0,0}\widehat{\bm{K}}_{N}\left(\bm{\theta}\right)}\big)$ is cheap because they only require computation of finite summations of terms with closed-form expressions. The only computational bottleneck is the matrix $\bm{A}\left(\bm{\theta}\right)$, which is a component of $\widehat{\bm{J}}_{N}\left(\bm{\theta}\right)$ and is defined as
@@
A.7 Gradient and Hessian expressions
@@
A.7.1 First order expressions
@@ A.7.2 Ancillary derivatives
A.7.3 Second order expressions
The (symmetric) Hessian matrix of the marginal density is given by
@@ A.7.3 Second order expressions
$\qquad\frac{1}{(1-\rho^{2})^{3/2}}\Bigg[\phi^{\prime}\left(\tau^{*}_{y}\left({x;\bm{\theta}}\right)\right)\left(\rho\tau_{y}-\frac{x-\mu}{\sigma}\right)\frac{\partial\tau^{*}_{y}\left({x;\bm{\theta}}\right)}{\partial\rho}+\tau_{y}\phi\left(\tau^{*}_{y}\left({x;\bm{\theta}}\right)\right){\color[rgb]{0,0,0}-}$
$\qquad\qquad\qquad\qquad\phi^{\prime}\left(\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)\right)\left(\rho\tau_{y-1}-\frac{x-\mu}{\sigma}\right)\frac{\partial\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)}{\partial\rho}{\color[rgb]{0,0,0}-}\tau_{y-1}\phi\left(\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)\right)\Bigg]$
@@
$\frac{\partial^{2}{p}_{Y|X}\left(y\ |\ x;\bm{\theta}\right)}{\partial{\rho}\partial{\mu}}$ $=\frac{1}{\sigma\sqrt{1-\rho^{2}}}\Bigg[\bigg(1+\frac{\rho^{{\color[rgb]{0,0,0}2}}}{1-\rho^{2}}\bigg)\bigg(\phi\left(\tau^{*}_{y}\left({x;\bm{\theta}}\right)\right)-\phi\left(\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)\right)\bigg)+$
@@
$\frac{\partial^{2}{p}_{Y|X}\left(y\ |\ x;\bm{\theta}\right)}{\partial{\rho}\partial{(\sigma^{2})}}$ $=\frac{x-\mu}{{\color[rgb]{0,0,0}2}\sigma^{3}\sqrt{1-\rho^{2}}}\Bigg[\bigg(1+\frac{\rho^{{\color[rgb]{0,0,0}2}}}{1-\rho^{2}}\bigg)\bigg(\phi\left(\tau^{*}_{y}\left({x;\bm{\theta}}\right)\right)-\phi\left(\tau^{*}_{y-1}\left({x;\bm{\theta}}\right)\right)\bigg)+$
@@
for $k\in\{1,\dots,r-1\}$, and, finally,
$\frac{\partial^{2}{p}_{Y|X}\left(y\ |\ x;\bm{\theta}\right)}{\partial{\tau_{i}}\partial{\tau_{j}}}=0$
for $i\neq j$.
We verified the correctness of all analytical expressions of derivatives provided in this Appendix A.7 using numerical approximations. The corresponding code is provided in the online replication material.
@@ Appendix B Algorithm for upper bound of raw weights
    Repeat Step 1 for all $y\in\mathcal{Y}$ and return the upper bound ${\color[rgb]{0,0,0}M}_{\alpha}(\bm{\theta})=\max\left\{S^{*}_{y}:y\in\mathcal{Y}\right\}$.
@@ C.3 Additional simulation results
Figure C.3: Fraction of the 5,000 simulation repetitions in which unconstrained optimization via BFGS did not converge, so constrained Nelder-Mead optimization was subsequently used instead.
@@
Figure C.4: Fraction of the 5,000 simulation repetitions in which a point estimator did not numerically converge—neither with the unconstrained BFGS nor the constrained Nelder-Mead algorithm—(top), and its associated standard error could not be computed due to singularity of $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator (bottom). The point estimators always converged, resulting in overplotted lines.
Recall from Section 6.4 that by default, our implementation first tries unconstrained optimization via the BFGS algorithm, and, in case of nonconvergence or an error, subsequently tries constrained optimization via the Nelder-Mead algorithm. Figure C.3 visualizes the frequency with which unconstrained optimization converged. In case of low contamination fractions, unconstrained optimization nearly always succeeded. Conversely, as the contamination fraction increases to higher values, unconstrained optimization fails more often, so a constrained algorithm must be used instead. Similar behavior was observed in all other simulations in this paper.
@@
Furthermore, Figure C.4 illustrates the fraction of the 5,000 repetitions in which the correlation estimate $\widehat{\rho}_{N}$ of given estimator $\widehat{\bm{\theta}}_{N}$ did not numerically converge to a solution, neither with the unconstrained BFGS nor the constrained Nelder-Mead algorithm (see Section 6.4). In similar fashion, it visualizes the fraction in which its associated estimated standard error $\widehat{\textnormal{SE}}\left(\widehat{\rho}_{N}\right)$ could not be computed due to singularity of matrix $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator in Section 6.2. The point estimates $\widehat{\rho}_{N}$ always numerically converged to a solution across all configurations and repetitions. Up to (and including) contamination fraction $\varepsilon=0.3$, the standard errors could virtually always be computed. Only in extremely high-contamination settings with $\varepsilon>0.3$, as the estimators gradually break down (Figure C.2), standard errors could occasionally also not be computed. In particular, at the highest considered contamination fraction $\varepsilon=0.49$, the estimator with tuning constant $\alpha=1$ fails to compute standard errors in about half of the repetitions. However, this estimator itself seems to be unstable in this setting to begin with (see Figure 2), which may cause frequent singularity of matrix $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$.
C.4 Discussion on robustifying the two-step approach
As described in Section 3.3, polyserial correlation is often estimated through the two-step approach in (7), which is computationally faster but less efficient than joint ML estimation. It turns out that the two-step approach is also a maximum likelihood (ML) estimator in which model parameters are estimated sequentially rather than jointly. In particular, it can be shown (Muthén,, 1984) that
  • the first-step estimators $\left(\widehat{\mu}_{\textrm{TS}},\widehat{\sigma}^{2}_{\textrm{TS}}\right)$ of $(\mu,\sigma^{2})$ are ML estimates of a marginal location-scale normality model for $X$;
  • the first-step threshold estimators $\widehat{\tau}_{k,\textrm{TS}}$ are ML estimates of thresholds $\tau_{k},k=1,\dots,r-1,$ in marginal latent normality model for $Y$,
  • the second-step correlation estimator of $\rho$ is a conditional ML estimator in a bivariate normality model, conditional on the first-step estimates.
It is therefore natural to ask whether the two-step approach can also be robustified by using DPD estimators. Indeed, it is possible to apply separate DPD estimators to separately estimate each of the three models above in a robust manner, owing to the generality of the results in Basu et al., (1998). However, there are substantial theoretical and computational downsides of such a robustified two-stage approach, as we shall explain in the following.
First, using DPD for jointly estimating all model parameters already incurs an efficiency loss, as discussed in detail in Section 6.3. Since using DPD in a two-stage approach requires fitting three separate models, the efficiency loss of such a robustified two-step approach would potentially be substantial due to using three estimation procedures with diminished efficiency. One of the attractive features of joint DPD estimation (as proposed in Section 5) is that robustness can be gained with only a relatively minor efficiency loss of less than 2% (see Section 6.3).
Second, every DPD estimator requires numerically solving an integral over the postulated model’s density to the power of $(1+\alpha)$ in every iteration of the employed optimization routine (Basu et al.,, 1998, Eq. 2.2). This numerical integral is the main computational bottleneck that causes the robust estimator in Section 5 to be computationally slower than joint ML. Recall that in our case, the integral is one-dimensional because the $Y$-dimension is a computationally cheap sum of $r$ summands; see Eq. (14). In a robustified two-step approach, however, there are three separate DPD estimators, each of which requires numerically solving a one-dimensional integral in every iteration. Hence, we expect the robustified two-stage approach to take even longer to compute than simultaneous estimation because the former requires separately solving three univariate integrals, while the latter only needs to solve one. While the main advantage of the two-step approach as in Olsson et al., (1982) is reduced computing time, such an advantage would therefore be absent (even reversed) in a DPD-based robustification thereof.
Thus, overall, robustifying the two-step approach via DPD would result in a procedure that is slower to compute and less efficient than joint estimation. Consequently, we believe that the latter approach is strictly preferable to the former.